OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b? 4, 1 -19, 40 20, 1 30, 9 -8, -29

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (solomonzelman):

you can write a system of equations, \(\large\color{black}{ a-b=21 }\) \(\large\color{black}{ 5a-2b=18 }\)

OpenStudy (solomonzelman):

I was thinking to put absolute values, but though it would in fact be wrong to do that.

OpenStudy (solomonzelman):

can you solve the system?

OpenStudy (skullpatrol):

Any questions @zxcvbnm08765 ?

OpenStudy (anonymous):

no @SolomonZelman

OpenStudy (solomonzelman):

lets start from multiplying the first equation of the system times -2, or times 5 (which ever way you like) and adding the equations together.

OpenStudy (solomonzelman):

I mean or times -5. not times 5.

OpenStudy (anonymous):

is a=-21 and b = -24/5

OpenStudy (solomonzelman):

you solved the system in your head?

OpenStudy (anonymous):

yea